OpenStudy (anonymous):

Which is the exact value of the expression sqrt{96}+ sqrt{150} - sqrt{54} Which is the exact value of the expression sqrt{96}+ sqrt{150} - sqrt{54} @Mathematics

OpenStudy (anonymous):

\[4\sqrt{6} + 5\sqrt{6} - 3\sqrt{6} = 6\sqrt{6}\]

OpenStudy (anonymous):

What is the simplified form of 18g^6h^6 --------- 12g^9h^4

OpenStudy (anonymous):

\[\frac{3h^{2}}{2g^{3}}\]

OpenStudy (anonymous):

What is the exat value of the expression \[5\sqrt{5}+3\sqrt{20}-6\sqrt{3}\]

OpenStudy (anonymous):

do it by urself